# Evolutionary Optimization of Physics-Informed Neural Networks: Evo-PINN Frontiers and Opportunities

### v3

@@ IV-A3 Memetic Algorithms Can Navigate Complex PINN Landscapes
There is a clear theoretical rationale supporting Evo-PINNs, yet there are only a handful of related works in the literature that explore such ideas. A barrier to the widespread use of EAs has been their high sample complexity, a consequence of updating and evaluating populations of solutions over multiple iterations. It is useful to consider hybridizing global evolutionary search with local gradient signals—thus inheriting the merits of both approaches. While the former can serve as an efficient basin-hopper for exploring rugged PINN landscapes, the latter facilitates rapid convergence to stationary points by exploiting local derivative information of loss functions.
There exists a long history of research in evolutionary computation studying synergies of this kind, with adaptive memetic algorithms forming a class of methods tailored for this purpose [143]. In the deep learning literature, gradient-guided evolutionary approaches have been designed for training neural networks [144, 145], with Cui et al. [146] proving a simple theorem that the performance of an evolutionary stochastic gradient descent method with elitist selection strategy never degrades with increasing iterations. However, to the best of our knowledge, there is no comparable body of work for PINN. Davi and Braga-Neto [129] were able to show that combining a swarm optimization algorithm (a population-based method like EAs) with gradient descent could produce an ensemble of PINN models that out-performed pure gradient-based learning. This result illuminates unexplored avenues for unifying SGD with gradient-free EAs as complementary algorithms within a single framework for finding globally optimum PINNs.
The crafting of synergistic optimization frameworks to meld gradient-based and gradient-free learning poses fundamental challenges, including determining interaction mechanisms that best harness the strengths of the two approaches. Key questions to address include balancing the intensity of SGD versus EAs, ascertaining when gradient descent should take over from evolution, or selecting an appropriate subset of the evolving population for fine-tuning with gradient descent (as doing so for the entire population would be too expensive). Notably, the subfield of memetic computation in computational intelligence—where ‘memes’ are loosely defined as computationally-encoded units of information (e.g., local gradient signals) combined with global evolutionary search for effective problem-solving—offers a wealth of theories and methodologies that directly tackle these issues [147, 148].
@@ IV-A4 Implementation Considerations and Hardware Acceleration
Employing EAs to directly optimize the neural network parameters introduces two primary challenges. First, EAs typically involve numerous hyperparameters (e.g., mutation rates, crossover rates, and population size) that require careful tuning. Fortunately, a substantial body of research has explored adaptive strategies for EA hyperparameter control [149], enabling EAs to better accommodate optimization problems of varying scales and characteristics. Among these, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [150] stands out as a representative algorithm, widely applied across machine learning domains. It has also been implemented in several EA-based optimization frameworks, such as EvoJAX [151], which facilitate efficient deployment in machine learning tasks. These integrated and adaptive EA implementations help reduce the burden of manual hyperparameter tuning in neuroevolution applications. Second, the large-scale nature of neural network parameter spaces presents a significant challenge, as EAs often exhibit reduced efficiency in high-dimensional optimization problems. Nevertheless, as highlighted in [152], this does not imply that larger neural networks necessarily underperform smaller ones when optimized using EAs. In fact, the global search capability of EAs and their robustness to local minima can possibly enable even smaller networks to achieve competitive predictive performance. Furthermore, several advanced strategies—such as sparse optimization [153] and cooperative coevolution [154]—have been proposed to improve the scalability of EAs in large-scale settings. These methods offer valuable insights and serve as potential tools for addressing high-dimensional optimization challenges in the context of Evo-PINNs.
Together with algorithm-centric research tracks, hardware-accelerated neuroevolution offers a promising path for EAs to substitute or be hybridized with SGD for PINN training. Sung et al. [155] provide preliminary results in this regard, demonstrating that neuroevolution with JAX [151] could lead to convergence trends that are orders of magnitude faster than standard implementations. The authors have made available codes for a set of five benchmark problems [^1] Python source codes for neuro-evolving PINN using EvoJAX are available at https://github.com/nicholassung97/Neuroevolution-of-PINNs., representative of real-world phenomena in classical mechanics, heat and mass transfer, fluid dynamics and wave propagation, with associated performance metrics and baseline results to support future research into the development of customized EAs for PINN training.
@@ IV-B Neural Architecture Search for Inducing Physics Biases
The choice of architectures and hyperparameters for PINNs has often been guided by experiential priors drawn from standard deep learning. However, a growing body of literature shows that accurate physics-informed learning may call for bespoke architectures. For example, contrary to popular belief, it was found that more hidden layers do not necessarily mean better performance and can sometimes be detrimental. An automated neural architecture search via a bilevel programming formulation revealed that, for certain governing differential equations, a shallow neural network with many neurons was more suitable for PINNs [156]. Similar intriguing insights regarding the choice of activation function have also been uncovered. While squashing functions (e.g., tanh and sigmoid) and variants of the rectified linear unit are the mainstay in the deep learning literature, the use of Fourier features (i.e., sine and cosine) has gained prominence in physics tasks [89, 91], specifically for their ability to overcome spectral bias and capture periodic or high frequency patterns (e.g., shock fronts) common in various physics applications [157, 158, 159].
Given the large number of design choices for network architectures and hyperparameters, manual experimentation to identify an optimal design for every physical problem is impractical. Notably, there exists a rich history of research in neuroevolution that automates such design choices through the joint evolution of neural network topologies along with their parameters [160, 161, 162]. One such example of the fully automatic design of PINNs by a genetic algorithm with specially tailored selection, crossover, and mutation operators was proposed in [53]. Each chromosome of this method encodes a learning algorithm, architecture of the PINN model, activation function in the hidden layers, and the output activation. In a related approach, an asynchronous genetic algorithm that alternates between model selection and optimization of physics-informed neural state space dynamics models was presented in [163]. Daw et al. [95] proposed an evolutionary sampling procedure to incrementally accumulate sample points in regions of high PDE residuals for mitigating training failures.
@@
Figure 7: Results extracted from [51] where empirical comparisons between SGD and evolution strategies are provided. A PINN simulating projectile motion under gravitational forces on Earth, Mars, or the moon is optimized by transfer natural evolution strategies—dubbed tNES. The convergence trends from SGD, xNES, and tNES for selected scenarios are displayed (the bold lines indicate the mean convergence path of multiple independent runs and the shaded areas indicate their 10th-90th interpercentile ranges). Individual solutions given by multiple SGD and tNES (transfer from a relevant prior, i.e., source distribution obtained by solving a different gravity $g$ or launch angle $a_{0}$) optimization runs for one particular scenario ($g$=1.6, $a_{0}$=45) are plotted and compared against the ground truth.
@@ IV-C1 Learning from Related Tasks
In a scientific study, it is common for a single set of PDEs to be evaluated under different physics scenarios (tasks), e.g., variations in PDE parameters, and IC/BCs. Moreover, for dynamical problems with causality, PINNs can generally be formulated to predict sequences of physical states, such as at different crack steps in a crack propagation study [175] or at different time steps in a time-varying dynamical problem [176, 177]. Such PINN models are sequentially learned using solutions from previous steps as IC/BCs. Relevant task solving experiences (i.e., optimizing PINN model parameters) can naturally accumulate over time. This allows transfer learning and optimization algorithms to exploit reusable knowledge from past episodes and jump-start the physics-informed learning, hence achieving rapid convergence for new target task [178]. The transfer learning of PINN from simpler tasks has also been shown to improve solution accuracy on challenging tasks, such as high-frequency and multi-scale problems, compared to learning the model from scratch [179].
@@ IV-C2 Adaptive Transfer
Adaptive transfer learning can be a particular strategy to address the issues without the need for explicit a priori identification of relevant sources. For example, an adaptive transfer physics-informed learning framework was proposed in [183] to achieve effective transfer, even when there is significant variation in the PDE parameters between the source and target tasks. The approach treats the PDE parameters as trainable, hence, they can be adaptively updated from the source to the target parameter values, creating a low-loss path in the parameter space to guide the optimization from the source to the target task. Throughout this process, the loss is consistently kept at a low level (similar to the concept of a minimum energy path connecting two minima [184]), ensuring that the PINN solution remains very close to the physical solution and thus guaranteeing the stability of the training process.
@@
Figure 8: Meta-learning PINN with MAML (left) versus Baldwinian evolution (right). In MAML, the initial PINN weights ${\bf\it W}$ are learned using gradient-based method, such that they can be quickly fine-tuned (physics-informed learning) on new tasks. As the task-specific physics-informed learning is limited to one or a few gradient descent updates during training, such limited fine-tuning is usually insufficient for a meta-learned PINN to converge. In Baldwinian evolution, the weight distribution in the pre-final nonlinear hidden layers ${\bf\it\tilde{w}}$ and learning hyperparameters $\lambda$ are jointly evolved. The task-specific physics-informed learning is performed on the output layer weights $w$ (segregated from ${\bf\it\tilde{w}}$) with a 1-step pseudoinverse operation at both training and test time (for linear ODE / PDEs).
@@ IV-D1 Meta-Learning in PINNs
Meta-learning algorithms provide a systematic framework for reusing knowledge from a distribution of learning tasks. This allows a meta-learned PINN model to quickly adapt to new physics-informed learning tasks with minimal training. Crucially, meta-learning offers a pathway towards a generalizable PINN over a broad class of physics scenarios, including different PDE families and geometries, as highlighted in Subsection III-B2. For example, the MAML and Reptile algorithms have been recently explored in the context of PINN [118, 119, 117]. They seek to meta-learn an optimal model initialization in the space of network weights from which the network can rapidly converge to a solution that minimizes the physics-informed loss of new tasks. This is achieved by a task-specific fine-tuning procedure during the meta-learning stage through an inner-outer loop training framework. This problem can be mathematically posed as a two-stage or bi-level optimization. The network weights are trained explicitly (in the outer loop) by optimizing the model performance over a batch of physics-informed learning tasks after one or a few gradient descent updates (at the inner loop) from the learned initialization.
@@ IV-D2 Neural Baldwinism
Figure 9: Schematic to illustrate new tasks arising from family of PDEs problem: S1 change to new PDE ($\gamma$ and $\delta$ parameters), source term $q$, and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. The solution for new PDE tasks (they are visually indistinguishable from the ground truth) can be quickly obtained in the order of milli-seconds by the Baldwinian-PINN [55]. In [55], CMA-ES was employed to evolve the Baldwinian-PINN models, demonstrating effective convergence in both physics residuals and MSE (108 training tasks). Superior performance is achieved with CMA-ES using initial standard deviations of 5 and 10. The mean MSE across 87 test tasks is below 5e-5.
@@
Recently, a meta-learning PINN framework based on the evolutionary principle of Baldwinism—dubbed Baldwinian PINN [55]—has been proposed to address the training challenges faced in conventional gradient-based meta-learning algorithms such as MAML. The Baldwin effect in nature explains the development of phenotypic plasticity among a population’s individuals [188], whereby their young are “born ready” with strong learning biases to perform a wide range of tasks. In order to pre-wire such learning ability into the initial connection strengths of a neural network, enabling it to be performant over a distribution of physics-informed learning tasks, an algorithmic realization of neural Baldwinism is examined in the context of PINNs. The algorithm tackles the two-stage optimization problem in meta-learning by coupling hidden layer evolution, for better generalization across tasks, with lifetime-learning at the output layer to specialize to particular tasks. Task-specific learning is limited to the output layer only [189, 190, 191], allowing the model to solve new linear PDE problems with a one-step pseudoinverse operation (or a few pseudoinverse steps for nonlinear PDE problems). The pseudoinverse operation has a closed form and is therefore fast. Unlike gradient-based MAML, the global evolutionary search in Baldwinian-PINNs is better equipped to handle the two-stage/bi-level problem formulation where exact gradients may be difficult to compute. Being a gradient-free algorithm, it does not require unrolling and differentiating through the computation graph, making Baldwinian-PINN an efficient method that scales well with the number of GPUs available. Fig. 8 illustrates the contrast between MAML and Baldwinian evolution for PINN [^2] Python source codes for Baldwinian-PINN are available at https://github.com/chiuph/Baldwinian-PINN..
@@
Experiments in [55] clearly illustrate how Baldwinian-PINN can consistently solve new physics scenarios within a diverse task distribution with orders of magnitude improvement in speed and accuracy compared to a single PINN model. These PDEs include both linear equations such as the linear convection-diffusion equation, Helmholtz equation and Poisson equation, and non-linear equations such as the Burger’s equation, Allen-Cahn equation and diffusion-reaction equation. Examples of the effective generalization of a Baldwinian-PINN to an entire PDE family with diverse output patterns is demonstrated in Fig. 9, on 2 task scenarios: S1 shows successful learning of solution for unseen set of PDEs, which include changes to PDE parameters, source term, and IC profile; while S2 shows effective extrapolation of solution to a longer time domain. Baldwinian-PINNs have clear benefits for the continual modeling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error. Crucially, Baldwinian evolution does not require explicit parameterization, and evolved Baldwinian-PINN can flexibly generalize to new ICs, BCs, and PDE source terms, a stark contrast to the need for existing parameterized PINNs to have pre-defined PDE, BC or IC parameters. Overall, the outstanding results in [55] compared to other meta-learning methods (Fig. 10) strongly motivate similar approaches in the spirit of Baldwinian-PINN.

### v5

@@ IV-A3 Memetic Algorithms Can Navigate Complex PINN Landscapes
There is a clear theoretical rationale supporting Evo-PINNs, yet there are only a handful of related works in the literature. A barrier to the widespread use of EAs has been their high sample complexity, a consequence of updating and evaluating populations of solutions over multiple iterations. It is useful to consider hybridizing global evolutionary search with local gradient signals—thus inheriting the merits of both approaches. While the former can serve as an efficient basin-hopper for exploring rugged PINN landscapes, the latter facilitates rapid convergence to stationary points by exploiting local derivative information of loss functions.
There exists a long history of research in evolutionary computation studying synergies of this kind, with adaptive memetic algorithms forming a class of methods tailored for this purpose [134]. In the deep learning literature, gradient-guided evolutionary approaches have been designed for training neural networks [135, 136], with Cui et al. [137] proving a simple theorem that the performance of an evolutionary stochastic gradient descent method with elitist selection strategy never degrades with increasing iterations. However, to the best of our knowledge, there is no comparable body of work for PINN. Davi and Braga-Neto [120] demonstrated that combining a swarm optimization algorithm (a population-based method like EAs) with gradient descent could produce an ensemble of PINN models that out-performed pure gradient-based learning. This result illuminates unexplored avenues for unifying SGD with gradient-free EAs as complementary algorithms within a single framework for finding globally optimum PINNs.
The crafting of synergistic optimization frameworks that meld gradient-based and gradient-free methods, however, remains challenging—particularly in identifying interaction mechanisms that effectively harness the strengths of both approaches. Key questions to address include balancing the intensity of SGD versus EAs, ascertaining when gradient descent should take over from evolution, or selecting an appropriate subset of the evolving population for fine-tuning with gradient descent (as doing so for the entire population would be too expensive). Notably, the subfield of memetic computation—where memes are loosely defined as computationally-encoded units of task-specific information (e.g., local gradient signals) combined with global evolutionary search—offers a wealth of theories and methodologies that tackle these issues [138, 139].
@@ IV-A4 Implementation Considerations and Hardware Acceleration
While population-based EAs offer search dynamics that complement point-based gradient descent, it is important to recognize that current algorithms may not yet provide an “off-the-shelf” solution for PINN optimization. Indeed, a recent study based on exploratory landscape analysis shows that the features of optimization benchmarks typically used in the development of EAs fundamentally differ from those of neural network training tasks [140]. Excelling at benchmarks does not, therefore, ensure effectiveness in neuroevolution. Although such in-depth analysis in the specific context of PINNs is so far unavailable, an empirical comparison of various evolution strategies (ES) by Sung et al. [141] shows that the performance of the classical CMA-ES [142] is statistically similar to or better than modern variants of natural ES algorithms [143] as well as SGD, across a range of PINN tasks representing real-world phenomena in classical mechanics, heat and mass transfer, fluid dynamics, and wave propagation. A compelling reason to consider ES is the induced Gaussian-smoothing of loss functions, biasing flat minima that often depict better generalization performance as per the minimum description length principle [144]. All algorithms tested in [141] were implemented in the JAX framework [145], providing orders of magnitude faster convergence than standard implementations. The source code has been made publicly available [^1] Python source codes for neuro-evolving PINN using EvoJAX are available at https://github.com/nicholassung97/Neuroevolution-of-PINNs. to facilitate deeper analysis and development of customized EAs for PINNs.
Not only is there a plethora of EAs to choose from, effective application of an EA to any problem also involves careful tuning of numerous hyperparameters (e.g., learning step-size, population size, crossover rates). Interestingly, the theory of information-geometric optimization (IGO) [119] provides a mathematical unification of probabilistic model-based EAs—including CMA-ES, natural ES, OpenAI-ES [146], or techniques that replace the Gaussian model with a deep generative model [147]—such that even a complete EA can be thought of as just a specific hyperparameter setting of the general IGO update rule. Empirical comparisons of optimizers suggest that more general update rules, if appropriately configured, should not underperform the ones they can approximate [148]. It is therefore fortunate that there exists a valuable body of literature on the automatic hyperparameter configuration of EAs [149] that could be leveraged for PINN optimization problems of varying scales and characteristics. Examples include but are not limited to adaptive population sizing [150] and reinforcement learning for online control of learning step-sizes [151] or crossover rates [149].
Another major challenge in the application of EAs is that they may scale poorly with the dimensionality of the optimized parameters. This number can indeed grow large for deep neural networks, and theoretical analysis suggests that the required number of optimization steps scales linearly with dimensions for general non-smooth functions [152]. However, OpenAI’s seminal paper on ES [146] prominently demonstrates that, in practice, these algorithms effectively exploit key informative dimensions of the optimization problem. This allows their performance to scale with the (low) intrinsic dimensionality of the task, rather than with the full dimensionality of the search space. This observation has been supported by recent work as well [5], positioning ES as a scalable option for training large PINNs given the availability of distributed hardware on which population evaluations can be parallelized. For tasks where the intrinsic dimensionality remains high, the latest methods of problem decomposition, cooperative coevolution, and memetic algorithms are worthy of exploration [153], though these have yet to be investigated in the context of PINNs.
@@ IV-B Neural Architecture Search for Inducing Physics Biases
The choice of architectures and hyperparameters for PINNs has often been guided by experiential priors drawn from standard deep learning. However, a growing body of literature shows that accurate physics-informed learning may call for bespoke architectures. For example, contrary to popular belief, more hidden layers do not necessarily mean better performance and can sometimes be detrimental. An automated neural architecture search via a bilevel programming formulation revealed that, for certain governing differential equations, a shallow neural network with many neurons was more suitable for PINNs [154]. Similar intriguing insights regarding the choice of activation function have also been uncovered. While squashing functions (e.g., tanh and sigmoid) and variants of the rectified linear unit are the mainstay in the deep learning literature, the use of Fourier features (i.e., sine and cosine) has gained prominence in physics tasks [81, 83], specifically for their ability to overcome spectral bias and capture periodic or high frequency patterns (e.g., shock fronts) common in various physics applications [155, 156, 157].
Given the large number of design choices for network architectures and hyperparameters, manual experimentation to identify an optimal design for every physical problem is impractical. Notably, there exists a rich history of research in neuroevolution that automates such design choices through the joint evolution of neural network topologies along with their parameters [158, 159]. One such example of the fully automatic design of PINNs by a genetic algorithm with specially tailored selection, crossover, and mutation operators was proposed in [50]. Each chromosome of this method encodes a learning algorithm, architecture of the PINN model, activation function in the hidden layers, and the output activation. In a related approach, an asynchronous genetic algorithm that alternates between model selection and optimization of physics-informed neural state space dynamics models was presented in [160]. Daw et al. [87] proposed an evolutionary sampling procedure to incrementally accumulate sample points in regions of high PDE residuals for mitigating training failures.
@@
Figure 7: A PINN simulating projectile motion under gravitational forces on Earth, Mars, or the moon is optimized by transfer natural evolution strategies—dubbed tNES. The empirical comparisons of convergence trends between SGD, xNES, and tNES for selected scenarios are provided (the bold lines indicate the mean convergence path of multiple independent runs and the shaded areas indicate their 10th-90th interpercentile ranges). Individual solutions given by multiple SGD and tNES (transfer from a relevant prior, i.e., source distribution obtained by solving a different gravity $g$ or launch angle $a_{0}$) optimization runs for one particular scenario (gg=1.6, $a_{0}$=45) are plotted and compared against the ground truth. Adapted with modifications from [48]. Original figure available at [arXiv:2101.01998].
@@ IV-C1 Learning from Related Tasks
In a scientific study, it is common for a single set of PDEs to be evaluated under different physics scenarios (tasks), e.g., variations in PDE parameters, and IC/BCs. Moreover, for dynamical problems with causality, PINNs can generally be formulated to predict sequences of physical states, such as at different crack steps in a crack propagation study [172] or at different time steps in a time-varying dynamical problem [173, 174]. Such PINN models are sequentially learned using solutions from previous steps as IC/BCs. Relevant task solving experiences (i.e., optimizing PINN model parameters) can naturally accumulate over time. This allows transfer learning and optimization algorithms to exploit reusable knowledge from past episodes and jump-start the physics-informed learning, hence achieving rapid convergence for new target task [175]. The transfer learning of PINN has also been shown to improve solution accuracy on challenging tasks, such as high-frequency and multi-scale problems, compared to learning the model from scratch [176].
@@ IV-C2 Adaptive Transfer
Adaptive transfer learning can be a particular strategy to address the issues without the need for explicit a priori identification of relevant sources. For example, an adaptive transfer physics-informed learning framework was proposed in [180] to achieve effective transfer, even when there is significant variation in the PDE parameters between the source and target tasks. The approach treats the PDE parameters as trainable, hence, they can be adaptively updated from the source to target parameter values, creating a low-loss path in the parameter space to guide the optimization from source to target tasks. Throughout this process, the loss is consistently kept at a low level (similar to the concept of a minimum energy path connecting two minima [181]), ensuring that the PINN solution remains very close to the physical solution and guaranteeing the stability of the training process.
@@
Figure 8: Meta-learning PINN with MAML (left) versus Baldwinian evolution (right). In MAML, the initial PINN weights are learned using gradient-based method, such that they can be quickly fine-tuned (physics-informed learning) on new tasks. The task-specific fine-tuning is limited to one or a few gradient descents during training, which is usually insufficient for a meta-learned PINN to converge. In Baldwinian evolution, the weight distribution in the pre-final nonlinear hidden layers $\bf\it\tilde{w}$} and learning hyperparameters $\lambda$ are jointly evolved. The task-specific fine-tuning is performed on the output layer weights $w$ (segregated from $\bf\it\tilde{w}$}) with a 1-step pseudoinverse operation at both training and test time (for linear ODE / PDEs). Adapted with modifications from [52]. Original figure available at [arXiv:2312.03243].
@@ IV-D1 Meta-Learning in PINNs
Meta-learning algorithms provide a systematic framework for reusing knowledge from a distribution of learning tasks, allowing models to quickly adapt to new physics-informed learning tasks with minimal training. Crucially, meta-learning offers a pathway towards a generalizable PINN over a broad class of physics scenarios, including different PDE families and geometries, as highlighted in Subsection III-B. For example, the MAML and Reptile algorithms have been recently explored in the context of PINN [109, 110, 108]. They seek to meta-learn an optimal model initialization in the space of network weights from which the PINN can rapidly converge to a solution that minimizes the physics-informed loss of new tasks. This is achieved by a task-specific fine-tuning procedure during the meta-learning stage through an inner-outer loop training framework. This problem can be mathematically posed as a two-stage or bi-level optimization. The network weights are trained explicitly (in the outer loop) by optimizing the model performance over a batch of physics-informed learning tasks after one or a few gradient descent updates (at the inner loop) from the learned initialization.
@@ IV-D2 Neural Baldwinism
Figure 9: Schematic to illustrate new tasks arising from family of PDEs problem: S1 change to new PDE ($\gamma$ and $\delta$ parameters), source term $q$, and IC profile for $t\in[0,2]$ (same time domain as train tasks), and S2 projection to longer time domain $t\in[0,4]$. The solution for new PDE tasks (they are visually indistinguishable from the ground truth) can be quickly obtained in the order of milli-seconds by the Baldwinian-PINN [52]. In [52], CMA-ES was employed to evolve the Baldwinian-PINN models, demonstrating effective convergence in both physics residuals and MSE (108 training tasks). Superior performance is achieved with CMA-ES using initial standard deviations of 5 and 10. The mean MSE across 87 test tasks is below 5e-5.Adapted with modifications from [52]. Original figure available at [arXiv:2312.03243].
@@
Recently, a meta-learning PINN framework based on the evolutionary principle of Baldwinism—dubbed Baldwinian PINN [52]—has been proposed to address the training challenges faced in conventional gradient-based meta-learning algorithms such as MAML. The Baldwin effect in nature explains the development of phenotypic plasticity among a population’s individuals [184], whereby their young are “born ready” with strong learning biases to perform a wide range of tasks. In order to pre-wire such learning ability into the initial connection strengths of a neural network, enabling it to be performant over a distribution of physics-informed learning tasks, an algorithmic realization of neural Baldwinism is examined in the context of PINNs. The algorithm tackles the two-stage optimization problem in meta-learning by coupling hidden layer evolution, for better generalization across tasks, with lifetime-learning at the output layer to specialize to particular tasks. Task-specific learning is limited to the output layer only [185, 186], allowing the model to solve new linear PDE problems with a one-step pseudoinverse operation (or a few pseudoinverse steps for nonlinear PDE problems). The pseudoinverse operation has a closed form and is therefore fast. Unlike gradient-based MAML, the global evolutionary search in Baldwinian-PINNs is better equipped to handle the two-stage/bi-level problem formulation where exact gradients may be difficult to compute. Being a gradient-free algorithm, it does not require unrolling and differentiating through the computation graph, making Baldwinian-PINN an efficient method that scales well with the number of GPUs available. Fig. 8 illustrates the contrast between MAML and Baldwinian evolution for PINN [^2] Python source codes for Baldwinian-PINN are available at https://github.com/chiuph/Baldwinian-PINN..
@@
Experiments in [52] clearly illustrate how Baldwinian-PINN can consistently solve new physics scenarios within a diverse task distribution with orders of magnitude improvement in speed and accuracy compared to a single PINN model. These PDEs include both linear equations such as the convection-diffusion, Helmholtz and Poisson equations, and non-linear equations such as the Burger’s, Allen-Cahn and diffusion-reaction equations. Examples of effective generalization of a Baldwinian-PINN to an entire PDE family with diverse output patterns is demonstrated in Fig. 9, on two task scenarios: S1 shows successful learning of solution for unseen set of PDEs, including changes to PDE parameters, source term, and IC profile; while S2 shows effective extrapolation of solution to a longer time domain. Baldwinian-PINNs have clear benefits for the continual modeling of dynamical systems through their versatility in handling ICs and ability to rapidly model and stitch time windows together with minimal error. Crucially, Baldwinian evolution does not require explicit parameterization, and evolved Baldwinian-PINN can flexibly generalize to new ICs, BCs, and PDE source terms, a stark contrast to the need for existing parameterized PINNs to have pre-defined PDE, BC or IC parameters. Overall, the outstanding results in [52] compared to other meta-learning methods (Fig. 10) strongly motivate similar approaches in the spirit of Baldwinian-PINN.